Two solutions of sodium hydroxide are in storage at a lab. One has a concentration of...

  1. Two solutions of sodium hydroxide are in storage at a lab. One has a concentration of hydrogen ions of 3.5x10-11, while the other has a concentration of 7.8x10-10. What are their pHs and which solution is more basic?
  1. You made a 0.3% KCl solution in Professor Forni’s lab but need to determine the Molarity of it for a paper. (MW of KCl is 74.55 g/mol)
Answer #1

The H+ is given in the question. It can be converted to pH using the formula=

H+= -logH+.

pH= - log (3.5*10^-11)

= 10.45

Similarly the pH of 7.8x10-10 = - log(7.8*10^-10)

= 9.1

The first solution is more basic., 10.45 is more basic than 9.1

15. The KCL used = 0.3%

= 0.3 gm in 100 ml

= 3 gm in 1000 ml

If you take 74.55 gm in 1000 ml it will be 1 mole solution. (Given in the question).

So, 3 gm is how many moles?

= (1*3)/74.55

= 0.04 moles

= 40 mmoles.

= 40 mM.